# Conditional existence of maximizers for the Tomas-Stein inequality for the sphere

### v1

@@ 1. Introduction
The Tomas-Stein’s inequality for the sphere or the adjoint Fourier restriction inequality for a compact subset [^1] The work of existence of extremizers for the Tomas-Stein inequality for the whole sphere is underway. of the sphere $\Gamma\subset S^{d}$ asserts that
@@ Acknowledgements
The first author would like to thank Michael Christ for many helpful discussions. This work grows out of an early discussion with him in 2010.
@@ 2. The refined Tomas-Stein estimate
To prove this lemma, we will use the Tao’s blinear restriction estimate for paraboloids [33]. Recently Frank, Lieb and Sabin directly prove this estimate in [14, Proposition 5.1] using similar ideas. We will provide a proof for completeness and follow the outline in Begout-Vargas [2]. Namely we first prove a Bourgain’s $X_{p}$ estimate. see e.g. [4, 23], and then deduce this estimate from it.
@@
We write the upper quarter sphere as $\Gamma\subset S^{d},\,\Gamma=\{(x,\phi(x)):\,|x|\leq\frac{1}{2},\phi(x)=\sqrt{1-|x|^{2}}\}$. For $f\in L^{2}(\Gamma,\sigma)$,
$\widehat{f\sigma}(t,x)=\int e^{ix\cdot\xi+it\phi(\xi)}f(\xi,\sqrt{1-|\xi|^{2}})\frac{d\xi}{\sqrt{1-|\xi|^{2}}}=\int e^{ix\cdot\xi+it\phi(\xi)}\frac{f(\xi,\sqrt{1-|\xi|^{2}})}{\sqrt{1-|\xi|^{2}}}d\xi.$
@@
By a simple argument using Theorem C, we obtain the following estimate.
@@
Indeed, we choose one large cube $\tau_{0}\subset B(0,1)$ that has comparable sizes to $\tau,\tau^{\prime}$ and contains $\tau,\tau^{\prime}$. If necessary, we rotate $\tau_{0}$ such that the sides of $\tau_{0}$ are parallel to the coordinates. Neither $\widehat{f\sigma}$ nor $\|f\|_{L^{p}(\Gamma,\sigma)}$ is not affected by this rotation; the same applies to $\widehat{g\sigma}$ and $\|g\|_{L^{p}(\Gamma,\sigma}$. This idea has also been implemented in [14, Theorem A. 1]. Noting that the side length of $\tau_{0}$ is $2^{-j}$, by scaling,
$\widehat{f\sigma}(t,x)=2^{-jd}\int_{Q}e^{i2^{-j}x\cdot\eta+it\phi(2^{-j}\eta)}f(2^{-j}\eta)d\eta,$
and
$\widehat{g\sigma}(t,x)=2^{-jd}\int_{Q^{\prime}}e^{i2^{-j}x\cdot\eta+it\phi(2^{-j}\eta)}f(2^{-j}\eta)d\eta$
where $Q,Q^{\prime}$ is the cube obtained from $\tau,\tau^{\prime}$ under this scaling, respectively. Note that $Q,Q^{\prime}$ are of distance $1$. Then we use Tao’s sharp bilinear restriction estimate for elliptic surfaces, Theorem C, from which Theorem D follows.
@@
This has previously appeared in [34]. Indeed, this Whitney type decomposition is possible because for given $a,b$, we consider the longest coordinate side-length $l_{max}$. There exists $k\in\mathbb{N}$ such that
$2^{-k}<l_{max}\leq 2^{-k+1}.$
Now we do the binary divisions of each side. Then the two points $a,b$ locate inside different $2^{-k-1}$ cubes which are not adjacent but have adjacent parents. Note that these two little cubes are separated at least by a thin rectangle array of length $2^{-k-1}$ that is orthogonal to the longest side, and hence are not adjacent. Note that this is different from the usual Whitney decomposition, see e.g. [29, Chapter 1, Theorem 3].
@@
$\widehat{f\sigma}(t,x)\widehat{f\sigma}(t,x)$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}e^{i(x\cdot\xi+t\phi(\xi))}e^{i(x\cdot\eta+t\phi(\eta))}f(\xi)g(\eta)d\xi d\eta$
@@
The $(n+1)$- dimensional Fourier support of $\widehat{f_{k}^{j}\sigma}$ is $\tilde{\tau_{k}^{j}}=\{(\phi(\xi),\xi):\,\xi\in\tau_{k}^{j}\}$. Hence the support of $\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}$ is contained in
$\tilde{\tau_{k}^{j}}+\tilde{\tau_{k^{\prime}}^{j}}=\{\bigl{(}\phi(\xi)+\phi(\xi^{\prime}):\,\xi+\xi^{\prime}\bigr{)}:\,\xi\in\tau_{k}^{j},\xi^{\prime}\in\tau_{k^{\prime}}^{j}\}.$
Let $\phi$ be the elliptic phase defined in Tao, Vargas, and Vega [34, Page 968] and Tao [33, Page 1381]. Then
$\phi(\xi)+\phi(\xi^{\prime})$ $=-\frac{1}{2}|\xi|^{2}-\frac{1}{2}|\xi^{\prime}|^{2}+\epsilon\bigl{(}f(\xi)+f(\xi^{\prime})\bigr{)}$
$=\bigl{(}-\frac{1}{2}+O(\epsilon)\bigr{)}\bigl{(}|\xi|^{2}+|\xi^{\prime}|^{2}\bigr{)}=\bigl{(}-\frac{1}{2}+O(\epsilon)\bigr{)}\bigl{(}|\xi+\xi^{\prime}|^{2}+|\xi-\xi^{\prime}|^{2}\bigr{)},$
where $O(\epsilon)$ denotes a number comparable to $C\epsilon$ for some constant $C>0$. Indeed, the use of the big $O$ notation here is harmless because, after given $j,k$, our goal is to counting the number of orthogonal pairs of cubes $\tau_{k^{\prime}}^{j}$ such that $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$ in the sense of the definition at the beginning of this proof, see Inequalities (9) and (11) below. We see that the sum of the two sets above is contained in
$H_{j,k}=\{(b,a)\in\mathbb{R}\times\mathbb{R}^{d}:\,|a-\frac{k}{2^{j-1}}|\leq\frac{C}{2^{j}},\frac{1}{4}2^{-2j}\leq b+(\frac{1}{2}-O(\epsilon))|a|^{2}\leq 4d2^{-2j},k\in\mathbb{Z}^{d}\},$
For each $j,k$, $H_{j,k}$ is a the vertical truncation of an intercepted paraboloid with a cylinder in $\mathbb{R}^{d+1}$. Note that
(9) $\sum_{j}\sum_{k}\sum_{k^{\prime}:\,\tau_{k}^{j}\sim\tau^{j}_{k^{\prime}}}\chi_{H_{j,k}}\leq C_{d},$
for some constant $C_{d}>0$. Indeed, for each $j,k$, there are $C2^{d}$-$k^{\prime}$ such that $\tau_{j,k^{\prime}}\sim\tau_{j,k}$. Let $M=\max\{2[\ln(d+1)],2\}$. We decompose each $\tau_{k}^{j}$ into dyadic subcubes of sidelength $2^{-j-M}$. Consequently we have a corresponding decomposition of $\tau_{k}^{j}\times\tau_{k^{\prime}}^{j}$ as follows. Set $\mathcal{D}$ the family of multi-indices $(m,m^{\prime},l)\in\mathbb{Z}^{d}\times\mathbb{Z}^{d}\times\mathbb{Z}$, so that there exists $\tau_{k}^{l-M}$ and $\tau_{k^{\prime}}^{l-M}$ with $\tau_{m}^{l}\subset\tau_{k}^{l-M}$, and $\tau^{l}_{m^{\prime}}\subset\tau_{k^{\prime}}^{l-M}$ and $\tau_{k}^{l-M}\sim\tau_{k^{\prime}}^{l-M}$ where j=lMj=l-M. Then
$1/2,1/2]^{d}\times[-1/2,1/2]^{d}\setminus\beta=\cup_{\mathcal{D}}\tau_{m}^{l}\times\tau_{m^{\prime}}^{l}.$
Hence
$\|\widehat{f\sigma}\|^{2}_{L^{2r}(\mathbb{R}^{d+1})}=\|\widehat{f\sigma}\widehat{f\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}=\|\sum_{\mathcal{D}}\widehat{f_{m}^{l}\sigma}\widehat{g_{m^{\prime}}^{l}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}.$
Note that if $(m,m^{\prime},l)\in\mathcal{D}$, then the distance between $\tau_{m}^{l}$ and $\tau_{m^{\prime}}^{l}$ is bigger than $2^{-l+M}\geq d2^{-l}$ and smaller than $4\sqrt{d}2^{-l+M}$. We claim that there are rectangles $R_{m,m^{\prime},l}$ and $c=c_{d}$ so that $\tilde{\tau}_{m}^{l}\times\tilde{\tau}_{m^{\prime}}^{l}\subset R_{m,m^{\prime},l}$ and $\sum_{\mathcal{D}}\chi_{(1+c)R_{m,m^{\prime},l}}\leq C_{d}$. We postpone the proof of this claim to the end of the proof. Assuming it holds, by Theorem E, since $r<2$, we have
@@
$\|\sum_{\mathcal{D}}\widehat{f_{m}^{l}\sigma}\widehat{f_{m^{\prime}}^{l}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C_{d}\left(\sum_{\mathcal{D}}\|\widehat{f_{m}^{l}\sigma}\widehat{f_{m^{\prime}}^{l}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}.$
@@
$\left(\sum_{\mathcal{D}}\|\widehat{f_{m}^{l}\sigma}\widehat{f_{m^{\prime}}^{l}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}$
$\leq C_{d,p}\left(\sum_{l}\sum_{m}\sum_{m^{\prime}:\,(m,m^{\prime},l)\in\mathcal{D}}2^{ldr\frac{2-p}{p}}\|f_{m}^{l}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{m^{\prime}}^{l}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
Now for each $(m,l)$ there are at most $4^{d}2^{Md}$ indices $m^{\prime}$ such that $(m,m^{\prime},l)\in\mathcal{D}$. Hence
$\left(\sum_{l}\sum_{m}\sum_{m^{\prime}:\,(m,m^{\prime},l)\in\mathcal{D}}2^{ldr\frac{2-p}{p}}\|f_{m}^{l}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{m^{\prime}}^{l}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
$\leq C_{d}\left(\sum_{l}\sum_{m}2^{ldr\frac{2-p}{p}}\|f_{m}^{l}\|^{2r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
We will still have to justify the claim. Assume, for the sake of simplicity that $\tau_{m}^{l}\times\tau_{m^{\prime}}^{l}$ is in the first quadrant of $\mathbb{R}^{d}$. Then $\tilde{\tau}_{m}^{l}+\tilde{\tau}_{m^{\prime}}^{l}$ is contained on a set
$H_{m,m^{\prime},l}$ $:=\{(a,b)\in\mathbb{R}^{d}\times\mathbb{R}:\,a=\frac{m+m^{\prime}}{2^{l}}+\nu,\nu=(\nu_{1},\cdots,\nu_{n}),0\leq\nu_{i}\leq 2^{-l+1},$
$\frac{1}{4}\times 2^{-2l+2M}\leq b+\bigl{(}\frac{1}{2}-O(\epsilon)\bigr{)}|a|^{2}\leq 16d2^{-2l+2M}\}.$
Consider the paraboloid defined by
$b+\bigl{(}\frac{1}{2}-O(\epsilon)\bigr{)}|a|^{2}=\frac{1}{4}\times 2^{-2l+2M}.$
Take $\Pi_{m,m^{\prime},l}$ to be the tangent plane to this paraboloid at the point $(a_{0},b_{0})$, with
$a_{0}=(m+m^{\prime})2^{-l},b_{0}=\bigl{(}-\frac{1}{2}+O(\epsilon)\bigr{)}|a_{0}|^{2}+\frac{1}{4}\times 2^{-2l+2M}$
and passing through that point. Also consider the point $(a_{1},b_{1})$ with $a_{1}=a_{0}+(2^{-l+1},\cdots,2^{-l+1})$ and $b_{1}=\bigl{(}-\frac{1}{2}+O(\epsilon)\bigr{)}|a_{1}|^{2}-16d\times 2^{-2l+2M}$. Then the rectangle $R_{m,m^{\prime},l}$ is defined as the only rectangle having a face contained in that hyperplane, and the points $(a_{0},b_{0})$, and $(a_{1},b_{1})$ as opposite vertices, as indicated in Figure 2.
Figure 2. The refined Whitney type decomposition.
Due to the convexity of paraboloids [18, page 44], it follows that
$H_{m,m^{\prime},l}\subset R_{m,m^{\prime},l}.$
Moreover, one can see that, for small $c=c_{d}$ because of the consideration of the inscribed and circumscribed cubes to a sphere in $\mathbb{R}^{d}$,
(10) $\begin{split}&(1+c)R_{m,m^{\prime},l}\subset\{(a,b):\,a=(m+m^{\prime})2^{-l}+\nu,\nu=(\nu_{1},\cdots,\nu_{d}),|\nu_{i}|\leq C_{d}2^{-l+1},\\ &C^{\prime}_{d}2^{-2l+2M}\leq b+\bigl{(}\frac{1}{2}-O(\epsilon)\bigr{)}|a|^{2}\leq C^{\prime\prime}_{d}2^{-2l+2M}\}.\end{split}$
for some constants $C_{d},C_{d}^{\prime},C_{d}^{\prime\prime}$. Indeed, $C_{d}$ is natural. For $C_{d}^{\prime}$ and $C_{d}^{\prime\prime}$, the equation for the tangent plane $(a,b)\in\Pi_{m,m^{\prime},l}$ is
$(1-O(\epsilon))a_{0}(a-a_{0})+(b-b_{0})=0.$
Then
$|b-b_{0}|=\left|(-1+O(\epsilon))a_{0}(a-a_{0})\right|\leq\frac{|m+m^{\prime}|}{2^{l}}2^{-l+1}\leq 4d2^{M}2^{-2l}\leq C2^{-2l+2M}$
for some absolute constant $C$, since the points are located in the square of side length $2^{M}$. Here $\frac{M}{2}\leq\ln(d+1)\leq M$ and so $d\sim 2^{M}.$ This is how the constants $C_{d}^{\prime}$ and $C_{d}^{\prime\prime}$ are chosen.
Therefore we have
(11) $\sum_{\mathcal{D}}\chi_{(1+c)R_{m,m^{\prime},l}}=\sum_{l}\sum_{m}\sum_{m^{\prime}:\,(m,m^{\prime},l)\in\mathcal{D}}\chi_{(1+c)R_{m,m^{\prime},l}}\leq C_{d}.$
Indeed, we see that $H_{m,m^{\prime},l}$ is contained in the rectangle $R_{m,m^{\prime},l}$; and $(1+c)R_{m,m^{\prime},l}$ is contained in a variant of the $H_{m,m^{\prime},l}$ as in (10). Then given $m,l$, we count how many $m^{\prime}$. First we know that there are $C_{d}$’s $k$. Then from the relation that $\tau_{k}^{l-M}\sim\tau_{k^{\prime}}^{l-M}$, there are $C_{d}$’s $k^{\prime}$. Then in $\tau_{k^{\prime}}^{l-M}$, there are $C_{d}$’s $m^{\prime}$. So in total there are $C_{d}$’s $m^{\prime}$ if given $m,l$. Now we finish the proof of Theorem 2.3. ∎
@@
(12) $\|\widehat{f\sigma}\|_{L^{2+\frac{4}{d}}}\leq C\left(\sup_{(j,k)\in\mathbb{Z}\times\mathbb{Z}^{d}}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\mu}\|f\|^{1-\mu p}_{L^{2}(\Gamma,\sigma)}.$

### v2

@@ 1. Introduction
The Tomas-Stein’s inequality for the sphere or the adjoint Fourier restriction inequality for a compact subset of the sphere $\Gamma\subset S^{d}$ asserts that
@@ Acknowledgements
The first author would like to thank Michael Christ for many helpful discussions. This work grows out of an early discussion with him in 2010. The authors have used the Chatgpt and overleaf to edit the figures in the paper.
@@ 2. The refined Tomas-Stein estimate
To prove this lemma, we will use the Tao’s blinear restriction estimate for paraboloids [33]. Recently Frank, Lieb and Sabin directly prove this estimate in [14, Proposition 5.1] using similar ideas. We will provide a proof for completeness, though the outline is essentially contained in Begout-Vargas [2]. Namely we first prove a Bourgain’s $X_{p}$ estimate, Theorem 2.3. see the precursors in [4, 23], and then deduce this estimate from it. We remark this $X_{p,q}$ estimate of the Strichartz norm of the sphere restriction operator $\widehat{f\sigma}$ is the first time of proof in the literature.
@@
We write the upper quarter sphere as $\Gamma\subset S^{d},\,\Gamma=\{(x,\sqrt{1-|x|^{2}}):\,|x|\leq\frac{1}{2}\}$. Let $\phi(x)=1-\sqrt{1-|x|^{2}}$. For $f\in L^{2}(\Gamma,\sigma)$,
$e^{-it}\widehat{f\sigma}(t,x)=\int e^{ix\cdot\xi+it\phi(\xi)}f(\xi,\sqrt{1-|\xi|^{2}})\frac{d\xi}{\sqrt{1-|\xi|^{2}}}=\int e^{ix\cdot\xi+it\phi(\xi)}\frac{f(\xi,\sqrt{1-|\xi|^{2}})}{\sqrt{1-|\xi|^{2}}}d\xi.$
@@
By a simple argument using Theorem C, Frank, Lieb and Sabin obtained the following theorem [14, Theorem A. 1].
@@
This has previously appeared in [34] and [16, Proposition 4.24]. Indeed, given $\xi\neq\xi^{\prime}$, there is a unique maximal pair $\xi\in\tau$ and $\xi^{\prime}\in\tau^{\prime}$ such that
$\tau\sim\tau^{\prime},\text{ and }|\tau|=|\tau^{\prime}|,\text{ and }\operatorname{dist}(\tau,\tau^{\prime})\geq\frac{1}{\sqrt{d}}\operatorname{diam}(\tau).$
Let $\mathcal{D}$ denote the collection of such pairs $\tau\sim\tau^{\prime}$ as $\xi\neq\xi^{\prime}$ varies over $\frac{1}{2},\frac{1}{2}]^{d}.$ According to this definition, we have
$\sum_{(\tau,\tau^{\prime})\in\mathcal{D}}\chi_{\tau}(\xi)\chi_{\tau^{\prime}}(\xi^{\prime})=1,\text{ for a.e. }(\xi,\xi^{\prime})\in[-\frac{1}{2},\frac{1}{2}]^{d}.$
Note that since $\tau$ and $\tau^{\prime}$ are maximal, $\operatorname{dist}(\tau,\tau^{\prime})\leq 4\operatorname{diam}(\tau).$ In addition, this shows that given $\tau$, there is a bounded number of $\tau^{\prime}$ such that $(\tau,\tau^{\prime})\in\mathcal{D}$. That is to say,
$\forall\tau,\#\{\tau^{\prime},\,\tau^{\prime}\sim\tau\}\lesssim 1.$
Note that this is different from the usual Whitney decomposition, see e.g. [29, Chapter 1, Theorem 3].
@@
$e^{-2it}\widehat{f\sigma}(t,x)\widehat{f\sigma}(t,x)$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}e^{i(x\cdot\xi+t\phi(\xi))}e^{i(x\cdot\eta+t\phi(\eta))}f(\xi)g(\eta)d\xi d\eta$
@@
[…]
@@
$\operatorname{supp}\mathcal{F}_{t,x}\left(\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\right)\subset\tilde{\tau_{k}^{j}}+\tilde{\tau_{k^{\prime}}^{j}}=\{\bigl(\phi(\xi)+\phi(\xi^{\prime}):\,\xi+\xi^{\prime}\bigr):\,\xi\in\tau_{k}^{j},\xi^{\prime}\in\tau_{k^{\prime}}^{j}\}.$
Here $\mathcal{F}_{t,x}(f)(\tau,\xi)=(2\pi)^{-\frac{d+1}{2}}\int_{\mathbb{R}^{d+1}}e^{i\xi\cdot x+it\tau}f(\tau,\xi)d\tau d\xi$. The $\phi$ is the elliptic phase defined in Tao, Vargas, and Vega [34, Page 968] and Tao [33, Page 1381]. We follow the approach outlined in [14, Lemma A.2] or [16, Proposition 4.24]. The Taylor expansion leads to the formula
$\phi(\xi)+\phi(\xi^{\prime})$ $=2\phi(\frac{\xi+\xi^{\prime}}{2})+a(\xi,\xi^{\prime})|\xi-\xi^{\prime}|^{2},$
$\phi(\frac{\xi+\xi^{\prime}}{2})$ $=\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})+\frac{1}{2}\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\xi+\xi^{\prime}-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))$
$+b(\eta+\eta^{\prime},c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))|\xi+\xi^{\prime}-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})|^{2},$
where $c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ denotes the center of the cube $\tau_{k}^{j}+\tau_{k^{\prime}}^{j}$, for two functions $a$ and $b$ satisfying
$\frac{3}{8}\leq a(\xi,\xi^{\prime})\leq-\frac{1}{8},\,-\frac{3}{16}\leq b(\xi+\xi^{\prime},c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))\leq-\frac{1}{16},$
assuming that $\epsilon$ in defining the elliptic phase in Tao, Vargas, and Vega [34, Page 968] and Tao [33, Page 1381] is sufficiently small. Further the space-time Fourier transform
$\operatorname{supp}\mathcal{F}_{t,x}\left(\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\right)\subset R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}),$
where
$R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}):=\left\{(b,a):\,a\in\tau_{k}^{j}+\tau_{k^{\prime}}^{j},\,-2\leq\frac{b-2\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(a-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j}))}{\operatorname{diam}(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})^{2}}\leq-\frac{1}{32}\right\}.$
Note that in computing these bounds forming the set, we have used
$0\leq|a-c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})|\leq\operatorname{diam}(\tau_{k}^{j}),\,\operatorname{diam}(\tau_{k}^{j})\leq|\xi-\xi^{\prime}|\leq 4\operatorname{diam}(\tau_{k}^{j}),\operatorname{diam}(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})=2\operatorname{diam}(\tau_{k}^{j}).$
Again by the Taylor expansion, we have
$\frac{3}{16}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq\phi(\frac{a}{2})-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot\frac{1}{2}(a-c(\tau_{k}^{j}+\tau_{k}^{j}))\leq-\frac{1}{16}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}.$
Therefore for $(b,a)\in R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$, noting that $0\leq|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|\leq\operatorname{diam}(\tau_{k}^{j})$,
$\begin{split}b-2\phi(\frac{a}{2})&=-2\left(\phi(\frac{a}{2})-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\right)\\ &+b+2\left(-\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})-\nabla\phi(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\cdot(\frac{c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}{2})\right).\end{split}$
Then
$8\operatorname{diam}(\tau_{j}^{k})^{2}+\frac{1}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq b-2\phi(\frac{a}{2})\leq-\frac{1}{8}\operatorname{diam}(\tau_{k}^{j})^{2}+\frac{3}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}.$
That is to say,
$8\operatorname{diam}(\tau_{j}^{k})^{2}\leq b-2\phi(\frac{a}{2})-\frac{1}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2},\text{ and }b-2\phi(\frac{a}{2})-\frac{3}{8}|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|^{2}\leq-\frac{1}{8}\operatorname{diam}(\tau_{k}^{j})^{2}.$
Therefore if two pair of close cubes $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$ and $\tau_{m}^{n}\sim\tau_{m^{\prime}}^{n}$ are such that $R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ and $R(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})$ intersect, they must have a similar diameter. The same holds for the dilates $(1+\alpha)R(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})$ for some small $\alpha$ by the same argument. Indeed, we can take $\alpha<0.01$. Lastly if the diameters are in a finite number, the cubes are also in a finite number since their centers satisfy
$|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-c(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})|\leq|c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})-a|+|c(\tau_{m}^{n}+\tau_{m^{\prime}}^{n})-a|\leq 2\operatorname{diam}(\tau_{j}^{k}).$
Then we see that
(9) $\sup_{b,a}\sum_{(\tau_{k}^{j},\tau_{k^{\prime}}^{j})\in\mathcal{D}}\chi_{(1+\alpha)R(c(\tau_{k}^{j}+\tau_{k^{\prime}}^{j})}(b,a)\leq C_{d},$
for some constant $C_{d}>0$. Then
$\|\widehat{f\sigma}\|^{2}_{L^{2r}(\mathbb{R}^{d+1})}=\|\widehat{f\sigma}\widehat{f\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}=\|\sum_{\mathcal{D}}\widehat{f_{k}^{j}\sigma}\widehat{g_{k^{\prime}}^{j}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}.$
By Theorem E, since $r<2$, we have
$\|\sum_{\mathcal{D}}\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|_{L^{r}(\mathbb{R}^{d+1})}\leq C_{d}\left(\sum_{\mathcal{D}}\|\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}.$
@@
$\left(\sum_{\mathcal{D}}\|\widehat{f_{k}^{j}\sigma}\widehat{f_{k^{\prime}}^{j}\sigma}\|^{r}_{L^{r}(\mathbb{R}^{d+1})}\right)^{1/r}$
$\leq C_{d,p}\left(\sum_{j}\sum_{k}\sum_{k^{\prime}}2^{jdr\frac{2-p}{p}}\|f_{j}^{k}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{k^{\prime}}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r},$
where $p,q,r$ are assumed in Theorem D. Now for each $(j,k)$ there are at most $2d$ indices $k^{\prime}$ such that $\tau_{k}^{j}\sim\tau_{k^{\prime}}^{j}$. Hence
$\left(\sum_{j}\sum_{k}\sum_{k^{\prime}}2^{ldr\frac{2-p}{p}}\|f_{k}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\|f_{k^{\prime}}^{j}\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
$\leq C_{d}\left(\sum_{j}\sum_{l}2^{jdr\frac{2-p}{p}}\|f_{k}^{j}\|^{2r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r}$
The last sum is recognized to be $\|f\|_{X_{p,q}}$ with $q=2r=\frac{2(d+2)}{d}$ and $\frac{1}{p^{\prime}}>\frac{d+3}{d+1}\frac{1}{q}$. Thus
$\|\widehat{f\sigma}\|_{L^{q}(\mathbb{R}^{d+1})}\leq\|f\|_{X_{p,q}}.$
Now we finish the proof of Theorem 2.3. ∎
@@
(10) $\|\widehat{f\sigma}\|_{L^{2+\frac{4}{d}}}\leq C\left(\sup_{(j,k)\in\mathbb{Z}\times\mathbb{Z}^{d}}|\mathcal{C}_{k}^{j}|^{\frac{p-2}{2}}\int_{\mathcal{C}_{k}^{j}}|f|^{p}\right)^{\mu}\|f\|^{1-\mu p}_{L^{2}(S^{d})}.$